Lightning produces a maximum air temperature on the order of 8.7 × 103 K, whereas a nuclear explosion produces a temperature on the order of 8.8 × 106 K. Use Wiens displacement law to calculate the wavelength of the thermally- produced photons radiated with greatest intensity by each of these sources. Select the part of the electromagnetic spectrum where you would expect each to radiate most strongly. (a) lightning Amax 3.33e-7 Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. nm It radiates most strongly in the ultraviolet ? part of the spectrum. (b) nuclear explosion max 3.29e-10 Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. pm It radiates most strongly in the gamma ray /x-raypart of the spectrum.
Answer #1

Wien's displacement law is Lmax = b / T, where Lmax is the most common wavelength emitted.

b = 2.898×10^?3

a. Lmax = (2.898×10^?3)/ (8.7*10^3) = 3.33e-7 m = 333 nm (ans)

b. Lmax = (2.898×10^?3)/ (8.8*10^6) = 3.29e-10 m = 329 pm (Ans)

[Your answers are correct. They just want it in other unit]